A bag of candy has 6 orange pieces, 8 green pieces, and 11 red pieces. What is the probability that the first piece of candy you draw at random out of the bag is red or green?
Answer by wuwei96815(245)

Total number of pieces is 6+8+11 = 25.
Total number of red and green candies is 11+8 = 19.

Probability that first piece drawn will be either red or green is 19/25 = 0.76 = 76%.

Does that look right?